# Asteroseismic Structure Inversions of Main-Sequence Solar-like Oscillators with Convective Cores

### v1

@@ 4.3 KIC 11807274
We now turn to the star with the most significant differences inferred by our inversions, KIC 11807274. We show in Figure 5 the frequency differences and frequency separation ratios for this star. Our reference model is in full agreement with the observed values of $T_{\rm{eff}}$ and [Fe/H]. The largest frequency differences are seen in the lower order quadrupole frequencies, which exhibit a glitch structure not reproduced in any model. To understand how sensitive our results are to these discrepant frequencies, we repeat both our modeling and inversion procedure excluding the lowest three quadrupole models. Our fitting procedure results in the same model as we found using the entire mode set. We show the averaging kernels and inversion results of the reduced mode set compared with the full mode set in Figure 5. Removing these modes results in slightly different averaging kernels, most notably for the highest target radius, where the maximum of kernel amplitude is shifted towards the center of the star. The differences inferred with these new averaging kernels are smaller than with the full mode set, however significant differences remain.
Figure 5: Results of the modeling and modeset testing of KIC 11807274. The top-left figure shows the relative frequency differences, after correcting for surface effects. The top-right figure shows the frequency separation ratios of the observed star and our reference model (computed using uncorrected model frequencies). The bottom right figure shows the original averaging kernels and those constructed when excluding the three lowest frequencies $l=2$ modes. The bottom left figure shows the results of inversions using each mode set.
@@ 4.4 Changes in Input Physics
We now explore several changes to the physics of our models as a first step to resolve the differences inferred by structure inversions. As a full study of all possible changes to all the stars is beyond the scope of this work, we focus only on the star with the most significant differences revealed by our inversions, KIC 11807274. The inversion results are shown in Figure 2. In Figure 6, we show the changes to the $\hat{u}$ profile that result from three different changes to the physics. The first change we present is a change to core boundary mixing. Instead of calculating the overshoot using exponential overmixing, where only the composition of the overshooting region is changed, this model uses a step convective penetration scheme described in Appendix A.1, where both the composition and the temperature gradient are changed in the overshoot region. For this change, we find new model parameters from a new grid created with the changed overshooting scheme. This change results in a slightly larger convective core, which causes the spike in the relative difference around $r/R\approx 0.05$. Otherwise, the main difference in the $\hat{u}$ profile is within the convective core, below the radius where our inversions probe. In the region where our inversions are sensitive, the change to $\hat{u}$ is in the correct direction according to our inversion results, but it is far too small to resolve the differences. In fact, the change is smaller than the uncertainties of our inversions. Several works (Lebreton & Goupil, 2012; Deal et al., 2023) have suggested that a large amount of convective penetration at the base of the outer convection zone may explain glitches observed in the frequency separation ratios of F-type stars. We have tested this perscription as well and find a change to the $\hat{u}$ at the base of the convective envelope $r/R\sim 0.8$, but at the radius where our inversions are sensitive the structure is very similar to the model without this additional mixing. Hence, this change is unable to resolve the differences inferred by our inversions.
@@
The other changes we examined deal with the transport of chemical elements. We tested the effects of including element diffusion and gravitational setting only as well as accounting for diffusion, settling, and radiative levitation. For both of these models we kept the same overshoot implementation and initial parameters as our original reference model, although we used the age that best fit the observations. In the model including only diffusion and settling we used the inlist parameters of the diffusion_smoothness test suite case in MESA. In the model including diffusion, settling, and radiative levitation we adopt the MESA settings of the A0 model in Campilho et al. (2022). In both of the new models we find the largest differences around the base of the convection zone where the transport processes have made the convection zone deeper. In the regions probed by our inversions, however, the changes are small enough to be within the uncertainties of our inversion results.
@@
Figure 6: Results of varying the several mixing processes for models of KIC 11807274. Each plot shows the relative difference in $\hat{u}$ between our original reference model and a model with different physics. We indicate 0 with a dashed horizontal line, the uncertainties of our inversion with dotted horizontal lines, and indicate the inferred difference at our middle target radius with the solid horizontal line. The vertical shading indicates the fractional radii where our inversions are sensitive. In all three panels, the spike around $r/R\approx 0.1$ is due to differences in the boundary of the convective core. The larger changes in the center and right panels around $r/R\approx 0.8$ are due to differences in the depth of the outer convection zone.
@@ 5 Conclusions
In this work, we have extended the analysis from Buchele et al. (2024) (B24) to stars with convective cores. We found best-fit models from a grid of tracks computed with MESA by fitting the observed frequencies, effective temperatures, and metallicities. Using each of these best-fit models, we performed structure inversions to compare the internal structure of the model to that of the star. These results, combined with the results of B24, show that our models reproduce the internal structure of around half of the stars examined. In cases where we find significant differences, we see an even split between models with dimensionless squared isothermal sound speed that is higher than the star and cases where it is too low. In contrast to the results of the stars with radiative cores, we did not find any significant correlations with the properties of our reference models. We tested several changes to the physics in our models and found that the resulting changes to the model structure are much smaller than necessary to reproduce the structure inferred by our inversions.
In both B24 and this work, the structure differences inferred by inversions remain unexplained. In future work, we plan to test potential modifications to the physical ingredients in our stellar modeling using structure inversions. With these changes we aim to consistently improve our models of the stars with significant differences without introducing discrepancies for the stars we currently model well.
@@ A.2 Modeling Information
In Table A.2, we provide the non-seismic constraints used in our modeling procedure. For all but three stars, we take our spectroscopic measurements from Furlan et al. (Table 9 of 2018) and adopt their suggested uncertainties of 100K and 0.1 dex for $T_{\rm{eff}}$ and [Fe/H], respectively. Two of our stars, KIC 434952 and KIC 5773345, are not analyzed by Furlan et al. (2018). In these cases, we use the values from Mathur et al. (2017). The values in Furlan et al. (2018) for KIC 9965715 were found to be discrepant from other literature values, so instead we use the measurements from Morel et al. (2021). To reduce the computation time when finding a best-fit model, we calculate $\chi^{2}_{\rm{fit}}$ only for models within $6\sigma$ of the observed effective temperature and metallicity and $10\sigma$ of the observed luminosity. We primarily use the FLAME luminosity value from Gaia DR3 (Gaia Collaboration et al., 2016; Creevey et al., 2023, 2023), although some stars are only available in Gaia DR2 (Gaia Collaboration et al., 2018), or not at all.
Table 4 provides the parameters of our reference model for each star and the star’s category based on our inversion results as defined in Section 4. We compare our stellar parameters to those reported in Silva Aguirre et al. (2017) in Figure 7. In general, our values of mass and radius fall within the spread of values in Silva Aguirre et al. (2017) without any clear biases. Our models tend to show higher initial hydrogen mass fraction and lower initial helium mass fraction. We attribute this to our choice not to include diffusion in our models.
@@
Star $M[\rm{M}_{\odot}]$ $Y_{\rm{initial}}$ $Z_{\rm{initial}}$ $\alpha_{\rm{mlt}}$ $f_{ov}$ $X_{c}$ $\chi^{2}_{\rm{fit}}$ Inversion Results Category
KIC 1435467 1.3540 0.2794 0.0209 2.2396 0.0231 0.2689 2.32 H
KIC 2837475 1.3297 0.2584 0.0123 2.1497 0.0243 0.4000 7.2 A
KIC 3456181 1.3165 0.2736 0.0104 2.2254 0.0504 0.2191 6.1 H
KIC 3632418 1.2773 0.2496 0.0116 1.9093 0.0219 0.1037 2.29 A
KIC 4349452 1.1394 0.2867 0.0166 2.0684 0.0344 0.3853 1.13 H
KIC 5184732 1.1683 0.3309 0.0399 2.2366 0.0124 0.1589 27.77 A
KIC 5773345 1.4914 0.2443 0.0254 2.0132 0.0464 0.2997 3.8 A
KIC 5866724 1.2668 0.2619 0.0235 2.1377 0.0250 0.3511 1.82 L
KIC 6225718 1.1553 0.2617 0.0147 2.3102 0.0392 0.4493 7.73 HL
KIC 6508366 1.4378 0.2824 0.0175 2.1715 0.0394 0.2403 5.88 L
KIC 6679371 1.5490 0.2421 0.0140 2.1485 0.0110 0.1463 4.38 A
KIC 7103006 1.4718 0.2541 0.0192 2.1252 0.0294 0.3029 1.49 H
KIC 7206837 1.2928 0.2711 0.0191 1.9863 0.0355 0.4079 2.27 A
KIC 7510397 1.3352 0.2438 0.0134 2.1562 0.0168 0.0838 5.46 A
KIC 7670943 1.2531 0.2456 0.0169 2.3065 0.0165 0.2911 1.62 A
KIC 7771282 1.2384 0.2417 0.0137 2.0294 0.0401 0.2714 1.85 L
KIC 7940546 1.3297 0.2584 0.0123 2.1497 0.0243 0.1338 7.97 H
KIC 8179536 1.2186 0.2677 0.0174 2.1525 0.0329 0.4253 3.1 H
KIC 8228742 1.2124 0.2791 0.0128 2.1034 0.0261 0.0278 3.35 A
KIC 8292840 1.1336 0.2461 0.0099 1.9458 0.0127 0.1372 2.28 H
KIC 8379927 1.2308 0.2483 0.0259 2.1166 0.0085 0.5646 6.51 L
KIC 8866102 1.2175 0.2505 0.0139 2.1005 0.0026 0.2313 2.06 A
KIC 9139151 1.1872 0.2655 0.0240 2.3548 0.0143 0.4026 4.92 L
KIC 9139163 1.3016 0.2601 0.0167 2.0525 0.0370 0.4446 5.65 LH
KIC 9206432 1.3915 0.2849 0.0245 2.0034 0.0062 0.5158 4.49 A
KIC 9353712 1.4165 0.2483 0.0149 2.0718 0.0390 0.2046 1.94 A
KIC 9414417 1.3359 0.2789 0.0135 2.3704 0.0200 0.1195 2.79 A
KIC 9592705 1.4472 0.3096 0.0262 2.0641 0.0135 0.0848 3.29 H
KIC 9812850 1.2289 0.2637 0.0106 2.1612 0.0590 0.2986 1.27 H
KIC 9965715 1.2133 0.3272 0.0273 1.9746 0.0147 0.4504 20.35 A
KIC 10068307 1.3505 0.2640 0.0151 2.2139 0.0354 0.0988 7.49 A
KIC 10162436 1.3365 0.2766 0.0149 2.1053 0.0276 0.1281 3.87 A
KIC 10454113 1.2531 0.2456 0.0169 2.3065 0.0165 0.4990 9.16 A
KIC 10644253 1.2308 0.2483 0.0259 2.1166 0.0085 0.6138 2.5 H
KIC 10666592 1.5095 0.2403 0.0193 2.0247 0.0130 0.2321 1.86 A
KIC 10730618 1.2735 0.2600 0.0110 2.2181 0.0435 0.2946 3.46 A
KIC 11081729 1.1776 0.2775 0.0126 1.9915 0.0186 0.3516 4.65 L
KIC 11253226 1.3502 0.2450 0.0129 2.0507 0.0410 0.4803 7.3 A
KIC 11807274 1.2015 0.2445 0.0120 2.1044 0.0203 0.1399 3.65 H
KIC 12009504 1.1244 0.2856 0.0140 2.1580 0.0232 0.2365 4.7 A
KIC 12069127 1.5569 0.2474 0.0173 2.0223 0.0103 0.0847 1.8 A
KIC 12258514 1.1489 0.2656 0.0127 2.1986 0.0307 0.0890 10.74 A
KIC 12317678 1.2723 0.2517 0.0077 1.9616 0.0089 0.0771 4.0 A

### v2

@@ 4.3 KIC 11807274
We now turn to the star with the most significant differences inferred by our inversions, KIC 11807274. We show in Figure 6 the frequency differences and frequency separation ratios for this star. This star also shows the glitches in the ratios discussed in the previous section. Our reference model is in full agreement with the observed values of $T_{\rm{eff}}$ and [Fe/H]. The largest frequency differences are seen in the lower order quadrupole frequencies, which exhibit a glitch structure not reproduced in any model. To understand how sensitive our results are to these discrepant frequencies, we repeat both our modeling and inversion procedure excluding the lowest three quadrupole models. Our fitting procedure results in the same model as we found using the entire mode set. We show the averaging kernels and inversion results of the reduced mode set compared with the full mode set in Figure 6. Removing these modes results in slightly different averaging kernels, most notably for the highest target radius, where the maximum of kernel amplitude is shifted towards the center of the star. The differences inferred with these new averaging kernels are smaller than with the full mode set, however significant differences remain.
Figure 6: Results of the modeling and modeset testing of KIC 11807274. The top-left figure shows the relative frequency differences, after correcting for surface effects. The top-right figure shows the frequency separation ratios of the observed star and our reference model. The bottom right figure shows the original averaging kernels and those constructed when excluding the three lowest frequencies $l=2$ modes. The bottom left figure shows the results of inversions using each mode set.
@@ 4.4 Changes in Input Physics
The fact that we find significant differences in many stars suggests that the physics commonly used in stellar modeling codes may need to be modified. A full study seeking to prove the accuracy of one set of physical choices over another will require extensive modeling of all stars for which structure inversions can be used. This modeling effort should test as many changes as possible to the modeling physics and explore the effect of fitting to different observables, such as the frequency separation ratios. While we think that such work is important to continue improving stellar models, it is beyond the scope of this work. Instead, here we present a few simple tests as examples of the changes to modeling physics that could be studied with inversions. For this we use the star where our inversions infer the most significant differences, KIC 11807274. In Figure 7, we show the changes to the $\hat{u}$ profile that result from three different changes to the physics. The first change we present is a change to core boundary mixing. Instead of calculating the overshoot using exponential overmixing, where only the composition of the overshooting region is changed, this model uses a step convective penetration (Zahn, 1991) scheme described in Appendix A.1, where both the composition and the temperature gradient are changed in the overshoot region. For this change, we find new model parameters from a new grid created with the changed overshooting scheme. This change results in a slightly larger convective core, which causes the spike in the relative difference around $r/R\approx 0.05$. Otherwise, the main difference in the $\hat{u}$ profile is within the convective core, below the radius where our inversions probe. In the region where our inversions are sensitive, the change to $\hat{u}$ is in the correct direction according to our inversion results, but it is far too small to resolve the differences. In fact, the change is smaller than the uncertainties of our inversions. Several works (Lebreton & Goupil, 2012; Deal et al., 2023) have suggested that a large amount of convective penetration at the base of the outer convection zone may explain glitches observed in the frequency separation ratios of F-type stars. We have tested this prescription as well and find a change to the $\hat{u}$ at the base of the convective envelope $r/R\sim 0.8$, but at the radius where our inversions are sensitive the structure is very similar to the model without this additional mixing. Hence, this change is unable to resolve the differences inferred by our inversions.
@@
The other changes we examined deal with the transport of chemical elements. We tested the effects of including element diffusion and gravitational setting only as well as accounting for diffusion, settling, and radiative levitation. In contrast to our test of convective penetration, for both of these models we kept the same overshoot implementation and initial parameters (mass, composition, overshoot and mixing length parameters) as our original reference model. The age of our new model is allowed to differ from the age of the original model. We choose the age along our new track which best fits the observations. In the model including only diffusion and settling we used the inlist parameters of the diffusion_smoothness test suite case in MESA. In the model including diffusion, settling, and radiative levitation we adopt the MESA settings of the A0 model in Campilho et al. (2022). In both of the new models we find the largest differences around the base of the convection zone where the transport processes have made the convection zone deeper. In the regions probed by our inversions, however, the changes are small enough to be within the uncertainties of our inversion results. Our choice to keep the initial parameters of the models constant between the different chemical mixing prescriptions represents the simplest possible test. In a full work seeking to fully resolve the inferred differences, these parameters should be inferred from a full grid as this change of physics is known to change the inferred mass, radius, and age of the star (e.g. Deal et al., 2020; Moedas et al., 2022, 2024), although how these changes affect the internal structure is not discussed in these works. For our purposes here, we seek only to provide some examples of the types of physical changes that can potentially be tested with structure inversions.
@@
Figure 7: Results of varying the several mixing processes for models of KIC 11807274. Each line shows the relative difference in $\hat{u}$ between our original reference model and a model with different physics. We indicate 0 with a dashed horizontal line and plot our inversion results for this star in black points. The vertical shading indicates the fractional radii where our inversions are sensitive. In all three cases, the spike around $r/R\approx 0.1$ is due to differences in the boundary of the convective core. The larger changes in the $\hat{u}$ profiles around $r/R\approx 0.8$ are due to differences in the depth of the outer convection zone.
@@ 5 Conclusions
In this work, we have extended the analysis from Buchele et al. (2024) (B24) to stars with convective cores. We found best-fit models from a grid of tracks computed with MESA by fitting the observed frequencies, effective temperatures, and metallicities. Using each of these best-fit models, we performed structure inversions to compare the internal structure of the model to that of the star. These results, combined with the results of B24, show that our models reproduce the internal structure of around half of the stars examined. In cases where we find significant differences, we see an even split between models with dimensionless squared isothermal sound speed that is higher than the star and cases where it is too low. In contrast to the results of the stars with radiative cores, we did not find any significant correlations with the properties of our reference models. We presented three models constructed with varying model physics as an example of the kinds of changes that could be tested in detail in future work using structure inversions. However, in our simple tests we found that the resulting changes to the model structure are much smaller than necessary to reproduce the structure inferred by our inversions.
In both B24 and this work, the structure differences inferred by inversions remain unexplained. Particularly in light of the upcoming PLATO mission (Rauer et al., 2014, 2024), it is important to continue improving our stellar models of these types of stars. In addition to both improved forward modeling and detailed analysis of the glitches in the observed frequencies and frequency separation ratios, structure inversions are an important part of ensuring that future models reproduce not only the global properties of the stars, but also their internal structure. To that end, in future work we plan to test potential modifications to the physical ingredients in our stellar modeling using structure inversions. With these changes, we aim to consistently improve our models of the stars with significant differences without introducing discrepancies for the stars we currently model well.
@@ A.2 Modeling Information
In Table A.2, we provide the non-seismic constraints used in our modeling procedure. For all but three stars, we take our spectroscopic measurements from Furlan et al. (2018, , Table 9) and adopt their suggested uncertainties of 100K and 0.1 dex for $T_{\rm{eff}}$ and [Fe/H], respectively. Two of our stars, KIC 434952 and KIC 5773345, are not analyzed by Furlan et al. (2018). In these cases, we use the values from Mathur et al. (2017). The values in Furlan et al. (2018) for KIC 9965715 were found to be discrepant from other literature values, so instead we use the measurements from Morel et al. (2021). To reduce the computation time when finding a best-fit model, we calculate $\chi^{2}_{\rm{fit}}$ only for models within $6\sigma$ of the observed effective temperature and metallicity and $10\sigma$ of the observed luminosity. We primarily use the FLAME luminosity value from Gaia DR3 (Gaia Collaboration et al., 2016; Creevey et al., 2023), although some stars are only available in Gaia DR2 (Gaia Collaboration et al., 2018), or not at all.
Table 4 provides the parameters of our reference model for each star and the star’s category based on our inversion results as defined in Section 4. Note that these are the parameters of our reference model and should not be treated as inferred values of the stellar parameters. We compare our model parameters to those reported in Silva Aguirre et al. (2017) in Figure 8. For the stars in our sample included in Silva Aguirre et al. (2015) but not in Silva Aguirre et al. (2017), we compare our reference model parameters in Figure 9. In general, our values of mass and radius fall within the spread of values in Silva Aguirre et al. (2015) and Silva Aguirre et al. (2017) without any clear biases. This is in contrast to the composition of our models which show a clear bias in favor of higher initial hydrogen mass fraction and lower initial helium mass fraction. We attribute this to our choice not to include diffusion in our models. Despite this bias, our values are still within the range of values predicted by the various pipelines used in Silva Aguirre et al. (2015) and Silva Aguirre et al. (2017).
To compare the quality of our fits, we obtain the surface-term corrected frequencies of the best fitting models found using the YMCM pipeline of Silva Aguirre et al. (2015, 2017). We obtained results for a set of models constructed both with and without diffusion. In Figure 10, we plot the distributions of $\chi^{2}_{\rm{fit}}$ values for the YMCM models with and without diffusion and the reference models used for our inversions. We find that our overall distribution is similar, with fewer outliers resulting from our modeling procedure.
@@
Star $M[\rm{M}_{\odot}]$ $Y_{\rm{initial}}$ $Z_{\rm{initial}}$ $\alpha_{\rm{mlt}}$ $f_{ov}$ $X_{c}$ Age [Gyr] $\chi^{2}_{\rm{fit}}$ Inversion Results Category
KIC 1435467 1.3540 0.2794 0.0209 2.2396 0.0231 0.2689 2.6253 2.32 H
KIC 2837475 1.3297 0.2584 0.0123 2.1497 0.0243 0.4000 2.0137 7.2 A
KIC 3456181 1.3165 0.2736 0.0104 2.2254 0.0504 0.2191 2.9891 6.1 H
KIC 3632418 1.2773 0.2496 0.0116 1.9093 0.0219 0.1037 3.6243 2.29 A
KIC 4349452 1.1394 0.2867 0.0166 2.0684 0.0344 0.3853 3.0303 1.13 H
KIC 5184732 1.1683 0.3309 0.0399 2.2366 0.0124 0.1589 4.3569 27.77 A
KIC 5773345 1.4914 0.2443 0.0254 2.0132 0.0464 0.2997 3.1090 3.8 A
KIC 5866724 1.2668 0.2619 0.0235 2.1377 0.0250 0.3511 3.2562 1.82 L
KIC 6225718 1.1553 0.2617 0.0147 2.3102 0.0392 0.4493 2.7897 7.73 HL
KIC 6508366 1.4378 0.2824 0.0175 2.1715 0.0394 0.2403 2.3956 5.88 L
KIC 6679371 1.5490 0.2421 0.0140 2.1485 0.0110 0.1463 1.9634 4.38 A
KIC 7103006 1.4718 0.2541 0.0192 2.1252 0.0294 0.3029 2.3479 1.49 H
KIC 7206837 1.2928 0.2711 0.0191 1.9863 0.0355 0.4079 2.6459 2.27 A
KIC 7510397 1.3352 0.2438 0.0134 2.1562 0.0168 0.0838 3.3145 5.46 A
KIC 7670943 1.2531 0.2456 0.0169 2.3065 0.0165 0.2911 3.3181 1.62 A
KIC 7771282 1.2384 0.2417 0.0137 2.0294 0.0401 0.2714 4.2713 1.85 L
KIC 7940546 1.3297 0.2584 0.0123 2.1497 0.0243 0.1338 3.0299 7.97 H
KIC 8179536 1.2186 0.2677 0.0174 2.1525 0.0329 0.4253 2.5597 3.1 H
KIC 8228742 1.2124 0.2791 0.0128 2.1034 0.0261 0.0278 3.9881 3.35 A
KIC 8292840 1.1336 0.2461 0.0099 1.9458 0.0127 0.1372 3.6114 2.28 H
KIC 8379927 1.2308 0.2483 0.0259 2.1166 0.0085 0.5646 1.3710 6.51 L
KIC 8866102 1.2175 0.2505 0.0139 2.1005 0.0026 0.2313 2.4654 2.06 A
KIC 9139151 1.1872 0.2655 0.0240 2.3548 0.0143 0.4026 2.3697 4.92 L
KIC 9139163 1.3016 0.2601 0.0167 2.0525 0.0370 0.4446 2.3699 5.65 LH
KIC 9206432 1.3915 0.2849 0.0245 2.0034 0.0062 0.5158 1.0102 4.49 A
KIC 9353712 1.4165 0.2483 0.0149 2.0718 0.0390 0.2046 3.0382 1.94 A
KIC 9414417 1.3359 0.2789 0.0135 2.3704 0.0200 0.1195 2.6318 2.79 A
KIC 9592705 1.4472 0.3096 0.0262 2.0641 0.0135 0.0848 2.3208 3.29 H
KIC 9812850 1.2289 0.2637 0.0106 2.1612 0.0590 0.2986 3.7670 1.27 H
KIC 9965715 1.2133 0.3272 0.0273 1.9746 0.0147 0.4504 1.5278 20.35 A
KIC 10068307 1.3505 0.2640 0.0151 2.2139 0.0354 0.0988 3.4016 7.49 A
KIC 10162436 1.3365 0.2766 0.0149 2.1053 0.0276 0.1281 3.0031 3.87 A
KIC 10454113 1.2531 0.2456 0.0169 2.3065 0.0165 0.4990 1.6690 9.16 A
KIC 10644253 1.2308 0.2483 0.0259 2.1166 0.0085 0.6138 0.9882 2.5 H
KIC 10666592 1.5095 0.2403 0.0193 2.0247 0.0130 0.2321 2.2907 1.86 A
KIC 10730618 1.2735 0.2600 0.0110 2.2181 0.0435 0.2946 3.1335 3.46 A
KIC 11081729 1.1776 0.2775 0.0126 1.9915 0.0186 0.3516 2.4767 4.65 L
KIC 11253226 1.3502 0.2450 0.0129 2.0507 0.0410 0.4803 2.0173 7.3 A
KIC 11807274 1.2015 0.2445 0.0120 2.1044 0.0203 0.1399 4.3142 3.65 H
KIC 12009504 1.1244 0.2856 0.0140 2.1580 0.0232 0.2365 3.8576 4.7 A
KIC 12069127 1.5569 0.2474 0.0173 2.0223 0.0103 0.0847 2.1788 1.8 A
KIC 12258514 1.1489 0.2656 0.0127 2.1986 0.0307 0.0890 4.9999 10.74 A
KIC 12317678 1.2723 0.2517 0.0077 1.9616 0.0089 0.0771 2.7866 4.0 A
@@
Figure 9: Comparison of our model parameters to several pipelines used in Silva Aguirre et al. (2015) The results of each pipeline are indicated with a dot and the uncertainties of that result a shaded region of the same color. The composition information is only available for the YMCM and AMP pipelines.
Figure 10: Distribution of $\chi^{2}_{\rm{fit}}$, as defined in Equation 3 for models fit using the YMCM pipeline of Silva Aguirre et al. (2015, 2017) and constructed with and without diffusion, as well as, the $\chi^{2}_{\rm{fit}}$ distribution of the models used in this work. Models with $\chi^{2}_{\rm{fit}}\geq 20$ have been collapsed into the final bin.